Veida007:

fine the value of cube 2cube times 3 cube

mhanifa:

It is confusing, can you attach a screenshot please?

Veida007:

mhanifa thereyou go

Veida007:

mhanifa pls solve it in process

mhanifa:

It is: \[\sqrt[3]{2^33^3 } = \sqrt[3]{6^3} = (6^3)^{1/3} = 6^{3*1/3} = 6^1\]

mhanifa:

You just need the very last step

Veida007:

thanks

Veida007:

mhanifa pls dm me